# A Review of Hydrogen-Enabled Resilience Enhancement for Multi-Energy Systems

### v2

@@ VI-B Multi-type Hydrogen Facilities
In order to deal with the worst-case cost under extreme events, Dong et al.[22] proposed a co-planning method for hydrogen-based microgrids and fuel-cell bus operation centers to decide the capacity, location, and quantity of HRS and FC electric bus considering stochastic line outages incurred by extreme weather events. Note that the proposed planning method is based on a nested C&CG algorithm, which can solve the formulated $\max-\min-\max$ tr-level optimization problem efficiently. In [111], Liu et al. investigated a hydrogen pipeline hardening and HESS allocation problem for electric-hydrogen network against extreme weather. Then, the above problem was formulated based on two-stage distributionally robust optimization framework, where the first stage focuses on pre-disaster decision under the given budget and the allowed hydrogen leakage risk, the second stage focuses on minimizing load shedding cost after events happen. The formulated problem was transformed into a two-stage MISOCP problem and was solved by C&CG problem. Simulation results indicated that the proposed strategy can minimize load shedding and control hydrogen pipeline leakage risk. Similar works could be observed in [112]. In [113], Wang et al. developed a resilient planning method for hybrid hydrogen and battery energy storage systems under extreme events based on particle swarm optimization and MILP, where particle swarm optimization is adopted to decide the system configuration (including installed location of fuel cell and battery; the rated power of fuel cell, battery, electrolyzer, and hydrogen tank; the capacity of battery) and the MILP related to the system operation cost minimization is solved by commercial solvers. In this work, parameter uncertainty and event uncertainty are not considered, while contingency modeling related to renewable energy reduction, load increase, and line capacity reduction is considered. Simulation results indicated that the hybrid energy system can improve the resilience index (i.e., load loss ratio) by 23.8% and 0.7%, compared with single battery energy storage or hydrogen energy storage, respectively.
Figure 10: The framework of the two-stage planning method
@@
Summary: To show the differences of existing works more clearly, their main features are summarized in Table VI and Table VII. It can be observed that not all works considered three kinds of uncertainties simultaneously, e.g., extreme event uncertainty, event-incurred contingency uncertainty, and system parameter uncertainty. Moreover, when more complex HMESs are considered (e.g., electricity-hydrogen-gas systems), the number of decisions related to different components becomes larger. Consequently, the formulated problems have more complex forms, e.g., MISOCP and MINLP. In addition, there are six kinds of modeling frameworks for planning enhancement problem of HMESs, i.e., single-level deterministic programming, bi-level programming, two-stage stochastic programming, three-level robust programming, three-level stochastic robust programming, and three-level distributionally robust optimization. Their modeling explanations, mathematical formulation, and solution method are summarized in Table 8. In row 1, $x$ denotes all decision variables, $c^{\top}x$ denotes the investment cost, $f(x)$ denotes other cost except investment cost; In row 2, $y$ denotes decision variables of lower-level optimization problem with objective function $f(x,y)$, $F(x,y)$ denotes the operation cost in the upper-level optimization problem; In row 3, $Q(x,\xi)$ denotes the operation cost under the scenario $\xi$; In row 4, $f(x,y,\xi)$ denotes the operation cost under the scenario $\xi$; In row 5, $f(x,y,u)$ denotes the operation cost under a decision $u$, which belongs to $\mathcal{U}(\xi)$; In row 6, $p$ denotes a probability distribution, $f(x,\xi)$ denotes the optimal operation cost under the scenario $\xi$ and decision $x$. Finally, the above works mainly focus on simultaneous line outages and/or device faults incurred by natural extreme events, while cascading faults and cyber attacks are seldom considered[114].
@@ VII-A Single-stage operation strategy
Figure 11: The key idea of the proposed scheduling method for the hydrogen-based multi-energy supply microgrid
@@
After the extreme events, some measures related to system topologies or resources could be adopted to restore load. For example, Zhu et al.[79] developed a resilience-oriented service restoration strategy for a power-hydrogen distribution system by taking some measures, e.g., network topology reconfiguration, scheduling of dispatchable generators and HESSs, and load prioritization. Similarly, Afsari et al.[74] proposed a service restoration strategy for microgrids with hydrogen storage based on network topology reconfiguration and the coordination of HESSs and hydrogen-based micro-gas turbines. Before network reconfiguration is performed, an HMES black-start is required. In [57], Lu et al. developed a black start strategy for hydrogen-integrated renewable grids with energy storage systems, where hydrogen fuel cells are used as black start resources. In addition to stationary HERs, MHERs could also be scheduled for restoration response. For example, Su et al.[80] proposed a restoration strategy for hydrogen-accommodated microgrids based on a two-stage stochastic programming framework, where the location allocation of mobile wind turbines is decided in the first stage according to the shortest-path information in the transportation network considering uncertainties in wind generation. Then, the expected costs of power outages will be minimized by scheduling mobile wind turbines and HESSs in microgrids jointly. In [76], Zhao et al. proposed a two-level restoration strategy for a distribution network with HRSs. As shown in Fig. 13, the distribution system operator schedules mobile energy sources (i.e., HTTs and HFCVs) to minimize the power imbalance among microgrids and reduce the unserved load in the upper level. Then, based on the allocation results, a microgrid energy management strategy is designed to minimize the system operation cost of each microgrid. Simulation results showed that the coordination of mobile energy sources and HFCVs can reduce the power imbalance and improve the self-healing ability of hydrogen-electricity systems. Different from above works, Wang et al.[120] proposed a restoration strategy for an integrated power and hydrogen distribution system with the consideration of power line and hydrogen pipeline outages caused by extreme events. When power outages happen, transmission line and hydrogen pipeline reconfiguration was first adopted to avoid islanding. Then, repair crews were dispatched to repair the fault lines and hydrogen pipelines. In addition, mobile battery-carried vehicles were dispatched to cooperate with repair crews to restore critical loads.
Figure 13: The framework of the two-level restoration strategy for a distribution network with HRSs
@@ VII-B Multi-stage operation strategy
Although the above single-stage operation strategies can enhance HMES resilience, they cannot implement the coordination between different stages. To this end, some multi-stage operation strategies for HMESs have been proposed. For example, Haggi et al. proposed a proactive rolling horizon optimization-based method for hydrogen systems to enhance the power system resilience[50], which can implement the seamless switching between normal and emergency modes. Under normal mode, hydrogen systems assist the power grid by acting as a load. Once the extreme event time is known based on the forecast, the hydrogen tank is filled to make preparation for emergency mode. In [86], Liu et al. developed a risk-averse receding horizon optimization method to support the proactive and emergency operation of the industrial park based on integrated hydrogen-electricity-heat microgrids, where the measure of proactive operation includes increasing the reserve capacity of FCs and that of emergency operation is to reschedule components for the electricity and heat load survivability. In [82], Wang et al. proposed a rolling dispatch method to enhance the resilience of hydrogen-penetrated Antarctic energy systems under extreme weather. The framework of the rolling dispatch method can be observed in Fig. 14, where three stages could be identified, i.e., normal scheduling, alert scheduling, and emergency scheduling. Moreover, the above-mentioned stages will transfer from one to another according to different event states. To be specific, when the extreme event information is known based on forecasting techniques at slot $t+s$, the normal scheduling stage with horizon length $h$ will transfer to the alert scheduling stage with rolling horizon length $h+\tau$. Here, $\tau$ is the extreme event duration. When the event beginning is detected, the alert scheduling stage will transfer to the emergency scheduling stage with a rolling horizon length $\tau$. Once the event ending is detected, the alert scheduling stage will transfer to the normal scheduling stage. Since the function of alert scheduling is similar to that of preventive response (i.e., making preparation for emergency response), their roles are regarded as equivalent in Table VI. Similar rolling horizon optimization-based methods considering economic operation, proactive preparation, and emergency operation can be observed in [89].
Figure 14: The framework of the rolling dispatch method with an alert mechanism
TABLE IX: Summary of Operation Enhancement for HMES (Single-stage)
Ref. Events Uncertainties Demand types Operation stages Problem types and solving methods Measure types
[85] Main grid power outages caused by extreme events $\times$ Electricity Preventive MILP; CPLEX Type-A
[84] Main outages caused by extreme events $\times$ Electricity Preventive MINLP; Generalized Benders Decomposition Type-A
[94] Line outages and generator disruption $\times$ Electricity, hydrogen Preventive MINLP Type-A
[78] Main grid power outages caused by extreme events Wind speed, solar radiation, active and reactive loads Electricity, hydrogen Preventive Stochastic MILP, fuzzy method Type-A
[29] Main grid power outages caused by natural disaster Active and reactive demands and renewable generation Electricity, hydrogen Preventive Stochastic MILP; CPLEX Type-A
[30] Main grid power outages caused by extreme events Electricity and hydrogen demands, renewable generators, temporal distances between HRSs Electricity, hydrogen Preventive Stochastic MILP; Gurobi Type-B
[51] Main grid power outages Renewable energy and electricity prices Electricity, hydrogen Emergency Stochastic MILP, CPLEX Type-D
[121] Main grid power outages caused by extreme events $\times$ Electricity, hydrogen Emergency Fuzzy multi-objective MISOCP; Gurobi Type-D
[117] Line outages and generator trips caused by ice disasters Line outage probability Electricity, hydrogen Emergency MILP, CPLEX Type-E
[31] Main grid power outage and insufficient natural gas supply caused by extreme events $\times$ Electricity, gas Emergency MILP; Alternating direction method of multipliers Type-E
[18] Component failures caused by hurricanes $\times$ Electricity, gas, heat Emergency MILP, Frank-Wolfe algorithm Type-F
[115] Component outage caused by extreme events Loads, renewables, electric vehicle charging behaviors Electricity, heat, cooling Emergency Stochastic MILP, GAMS solver Type-H
[119] Extreme PV and wind turbine generation PV and wind turbine generation Electricity, heat, hydrogen Emergency Stochastic MILP, Gurobi/CPLEX Type-E, Type-G
[59] False data injection cyber-attacks on electricity and hydrogen demand $\times$ Electricity, heat, hydrogen Emergency MILP; CPLEX Type-G
[79] Distribution line and pipeline outages caused by extreme events $\times$ Electricity, heat, hydrogen Restoration MISOCP; Relaxation inducement Type-K
[74] Line outages caused by natural disasters Load demand Electricity Restoration Chance-constrained adaptive distributionally robust MILP; CPLEX Type-K
[57] Large-scale blackouts caused by extreme events $\times$ Electricity Restoration MILP Type-L
[80] Line outages caused by extreme weather events Wind forecast Electricity Restoration Stochastic MILP; Gurobi Type-I
[76] Line outages caused by extreme events Renewable outputs, electric and thermal loads Electricity, heat, hydrogen Restoration Stochastic MILP; Gurobi Type-J
[120] Power line and hydrogen pipeline outages caused by extreme events $\times$ Electricity, hydrogen Restoration MILP, Gurobi Type-M
TABLE X: Summary of Operation Enhancement for HMES (Multi-stage)
@@
Ref. Events Uncertainties Demand types Operation stages Problem types and solving methods Hydrogen-related measure types
@@
[86] Critical component contingencies caused by extreme events Network contingencies Electricity, heat Preventive, emergency Stochastic MISOCP; Receding horizon method Type-A, Type-D
@@
[82] Outage of renewable energy units by storms Renewable energy Electricity, heat Preventive, emergency MILP; Gurobi Type-A, Type-D
@@
[89] Main grid power outages and wind power plant contingencies caused by hurricanes Extreme event starting time Electricity, heat, hydrogen Preventive, emergency Rolling-horizon MILP; Gurobi Type-A, Type-D
@@
[72] Line outages caused by hurricane Failed transmission lines, their failure, and repair times Electricity, hydrogen Preventive, emergency Stochastic multi-objective MILP Type-D
[23] Line outages and reduced PV generation caused by freezing rain diaster Line outages Electricity, heating, cooling, hydrogen Preventive, emergency MILP; Gurobi Type-A, Type-D
@@
[11] Line outages caused by extreme weather Line faults Electricity Preventive, emergency Two stage robust optimization; C&CG Type-B, Type-F
[73] Line outages caused by hurricane Damaged transmission lines, hurricane time, and repair time Electricity, heat, hydrogen Preventive, emergency Stochastic MILP Type-C, Type-D
[77] Main grid disconnection and line outages caused by extreme events Line faults Electricity Preventive, emergency Tri-level MILP; Nested C&CG Type-C, Type-F
@@
[81] Main grid power outages caused by extreme events Load, renewable energy, and travelling time Electricity, heat Preventive, restoration Stochastic MILP; Gurobi Type-D
@@
[87] Lines/equipment failures incurred by hurricanes Lines/component failure probability Electricity Preventive, restoration Stochastic multi-objective programming; NSGA-II Type-C, Type-J
@@
[75] Line outages caused by typhoons $\times$ Electricity, hydrogen Preventive, restoration MISOCP; CPLEX Type-A, Type-J, Type-K
[122] Main grid disconnection during emergencies $\times$ Electricity, heat, cooling, hydrogen Emergency, restoration MILP; CPLEX Type-G, Type-J
@@
[21] Line outage caused by extreme events Renewable generation, load demands, and charging level of EVs Electricity, heat, cooling, hydrogen Emergency, restoration Bi-level MILP; Adaptive ADMM algorithm Type-G, Type-J
@@
[123] Line outages and insufficient gas supply caused by extreme events Demand and renewable generation Electricity, heat, gas, hydrogen Emergency, restoration Stochastic MIQCP; Augmented ADMM algorithm Type-F, Type-D, Type-J
@@
TABLE XI: Representative Hydrogen-related Operation Measures for HMES resilience enhancement
@@
Preventive Type-A HESS pre-charging Increasing HESS level for reserve[50][75][82][84, 85, 86][23]
Type-B HTT scheduling HRS-based microgrid formation assisted by transport-based measures (e.g., HTTs and mobile hydrogen trucks) [30], pre-disaster HTT scheduling[11]
@@
Emergency Type-D HESS discharging Hydrogen tank-driven fuel cell for independent power supply during outages[51], discharging batteries and HESS[89], HESS and CHP work at maximum power[82], discharging HESS for serving HFCVs[23], discharging HESS for power supply[86]
Type-E Natural gas pipeline-assisted hydrogen use, hydrogen pipeline scheduling Utilizing hydrogen enriched natural gas for fuel support[31], using pipelines as large-scale energy storage[119], using hydrogen stored in the pipeline for power support[117]
Type-F HTT/hydrogen truck scheduling Using mobile hydrogen trucks for critical load support[18][19], fuel cell power supply supported by hydrogen from HTTs[11], dynamic re-routing of HTTs[77]
@@
Restoration Type-I Mobile hydrogen generation and storage Combined operation of mobile wind turbines and HESSs[80]
Type-J HFCV or hydrogen truck scheduling HFCV routing and dispatch[75][76][87], HFCV for power support[75], hydrogen fuel cell trucks are scheduled to HRSs for charging and dispatched to connection points for power supply[21][122]
Type-K HESS scheduling Scheduling HESS with other components[81], HESS discharging[74][75]

### v3

@@ VI-B Multi-type Hydrogen Facilities
In order to deal with the worst-case cost under extreme events, Dong et al.[22] proposed a co-planning method for hydrogen-based microgrids and fuel-cell bus operation centers to decide the capacity, location, and quantity of HRS and FC electric bus considering stochastic line outages incurred by extreme weather events. Note that the proposed planning method is based on a nested C&CG algorithm, which can solve the formulated $\max-\min-\max$ tr-level optimization problem efficiently. In [112], Liu et al. investigated a hydrogen pipeline hardening and HESS allocation problem for electric-hydrogen network against extreme weather. Then, the above problem was formulated based on two-stage distributionally robust optimization framework, where the first stage focuses on pre-disaster decision under the given budget and the allowed hydrogen leakage risk, the second stage focuses on minimizing load shedding cost after events happen. The formulated problem was transformed into a two-stage MISOCP problem and was solved by C&CG problem. Simulation results indicated that the proposed strategy can minimize load shedding and control hydrogen pipeline leakage risk. Similar works could be observed in [113]. In [114], Wang et al. developed a resilient planning method for hybrid hydrogen and battery energy storage systems under extreme events based on particle swarm optimization and MILP, where particle swarm optimization is adopted to decide the system configuration (including installed location of fuel cells and batteries; the rated power of fuel cells, batteries, electrolyzers, and hydrogen tanks; the capacity of batteries) and the MILP related to the system operation cost minimization is solved by commercial solvers. In this work, parameter uncertainty and event uncertainty are not considered, while contingency modeling related to renewable energy reduction, load increase, and line capacity reduction is considered. Simulation results indicated that the hybrid energy system can improve the resilience index (i.e., load loss ratio) by 23.8% and 0.7%, compared with single battery energy storage or hydrogen energy storage, respectively.
Figure 10: The framework of the two-stage planning method[100]
@@
Summary: To show the differences of existing works more clearly, their main features are summarized in Table VI and Table VII. It can be observed that there are typically three kinds of uncertainty processing methods in existing works. To better understand the differences among these methods, we provide a clear classification of them, as shown in Table VIII. Among them, scenario-based methods are suitable for explicitly modeling the evolution of uncertain events, and related planning problems typically follow the framework of stochastic programming. Instead of representing uncertainty by a finite number of scenarios, uncertainty-set-based methods adopt deterministic sets to describe uncertainties, e.g., interval sets and cardinality-constrained uncertainty sets. These methods do not require exact probability distributions, and related planning problems typically follow the robust optimization or stochastic robust optimization framework. When the exact probability distribution is unknown but partial statistical information is available, ambiguity-set-based methods can be adopted, which model uncertainty through a family of probability distributions. In this case, related planning problems typically follow the framework of distributionally robust optimization. Moreover, when more complex HMESs are considered (e.g., electricity-hydrogen-gas systems), the number of decisions related to different components becomes larger and more nonlinear or nonconvex constraints incurred by hydrogen equipment (e.g., electrolyzers, fuel cells, hydrogen pipelines, hydrogen compressors/liquefiers, and hydrogen storage tanks) are introduced. Consequently, the formulated problems have more complex forms (e.g., a MINLP problem is incurred due to the nonlinear hydrogen pipeline hydraulics (i.e., equation (28) in [109])) and advanced solution methods are expected. Furthermore, there are six kinds of modeling frameworks for planning enhancement problem of HMESs and we summarize their differences in Table IX. In row 1, $x$ denotes all decision variables, $c^{\top}x$ denotes the investment cost, $f(x)$ denotes other cost except investment cost; In row 2, $y$ denotes decision variables of lower-level optimization problem with objective function $f(x,y)$, $F(x,y)$ denotes the operation cost in the upper-level optimization problem; In row 3, $Q(x,\xi)$ denotes the operation cost under the scenario $\xi$; In row 4, $f(x,y,\xi)$ denotes the operation cost under the scenario $\xi$; In row 5, $f(x,y,u)$ denotes the operation cost under a decision $u$, which belongs to $\mathcal{U}(\xi)$; In row 6, $p$ denotes a probability distribution, $f(x,\xi)$ denotes the optimal operation cost under the scenario $\xi$ and decision $x$. Finally, the above works mainly focus on simultaneous line outages and/or device faults incurred by natural extreme events, while cascading faults and cyber attacks are seldom considered[115]. In addition, a few of them considered scenario generation for extreme events based on generative adversarial networks[47] and natural disaster models[105]. Although such methods have some advantages, they face some limitations in aspects of physical implausibility and bias correction. To show the advantages and disadvantages of other scenario generation methods, we provide a classification of extreme-event scenario generation methods in Table X, where representative methods, key idea, advantages, and limitations could be observed.
TABLE X: Classification of extreme-event scenario generation methods considering the uncertainty of extreme events themselves
Category Representative methods Key idea Advantages Limitations
Statistical sampling-based scenario generation Historical event selection[116]; bootstrap resampling[117]; block bootstrapping[118]; parametric distribution fitting with random sampling[47]; representative-period extraction[119] Generate extreme-event scenarios directly from historical observations or their statistical properties, such as empirical frequencies, fitted marginal distributions, or resampling rules Simple and transparent; easy to implement; low computational burden; easy to integrate into stochastic planning models; preserves some historical temporal characteristics Difficult to generate unprecedented or very rare events; limited ability to represent compound extremes and high-dimensional dependence
Dependence-aware probabilistic modeling Copula-based models[120][121]; multivariate parametric distributions[122]; Bayesian networks[123]; spatio-temporal probability modeling[124] Construct explicit joint probabilistic relationships among multiple variables, so that generated scenarios preserve inter-variable, spatial, and temporal dependence structures Better captures spatial correlation and temporal clustering; more suitable than simple resampling for multi-variable extreme-event modeling Model estimation becomes difficult in high-dimensional settings; tail dependence is hard to identify under limited extreme samples
Climate- and weather-model-driven scenario generation Regional climate models[105]; global climate model outputs[117]; numerical weather prediction models[125]; bias-corrected climate projections[126]; weather generators[127] Generate extreme-event scenarios from physically based simulation or projection of meteorological and climate processes, rather than only from historical observations Physically grounded; suitable for future climate conditions and long-return-period events; capable of representing persistent and evolving meteorological extremes Requires downscaling, bias correction, and large computational effort; outputs may contain structural uncertainty from the underlying climate/weather models
Artificial intelligence-enabled synthetic scenario generation Generative adversarial networks[47][128][125]; variational autoencoders[120]; diffusion models[129][130]; hybrid deep generative models[131] Learn complex nonlinear spatio-temporal patterns from data and generate synthetic extreme-event trajectories beyond direct historical resampling Flexible in handling high-dimensional and nonlinear patterns; potentially better at capturing complex spatio-temporal structures than classical approaches Usually data-hungry; may generate physically implausible scenarios without domain constraints; interpretability is weak
@@ VII-A Single-stage operation strategy
Figure 11: The key idea of the proposed scheduling method for the hydrogen-based multi-energy supply microgrid[18]
@@
TABLE XI: Summary of Operation Enhancement for HMES (Single-stage)
Ref. Parameter/event uncertainty Contingency uncertainty Uncertainty processing methods Operation stages Problem types and solving methods Measure types
[86] $\times$ Deterministic grid outage $\times$ Preventive MILP; CPLEX A
[85] $\times$ Deterministic grid outage $\times$ Preventive MINLP; Generalized benders decomposition A
[95] $\times$ Deterministic line and generator outages $\times$ Preventive MINLP A
[79] Wind speed, solar radiation, active and reactive loads Deterministic grid outage Scenarios Preventive Stochastic MILP, fuzzy method A
[29] Demands and renewable generation Deterministic grid outage Scenarios Preventive Stochastic MILP; CPLEX A
[30] Demands, renewable generators, HRS distances Deterministic grid outage Scenarios Preventive Stochastic MILP; Gurobi B
[51] Renewable energy and electricity prices Deterministic grid outage Scenarios Emergency Stochastic MILP; CPLEX D
[137] $\times$ Deterministic grid outage $\times$ Emergency Fuzzy multi-objective MISOCP; Gurobi D
[134] $\times$ Deterministic line outages and generator trips $\times$ Emergency MILP, CPLEX E
[31] $\times$ Deterministic grid outage and gas shortage $\times$ Emergency MILP; Alternating direction method of multipliers E
[18] $\times$ Deterministic component failures $\times$ Emergency MILP, Frank-Wolfe algorithm F
[76] Loads, renewables, electric vehicle charging behaviors Deterministic component outages Scenarios Emergency Stochastic MILP, GAMS solver H
[136] PV and wind turbine generation Extreme PV and wind turbine generation Scenarios Emergency Stochastic MILP, Gurobi/CPLEX E, G
[59] $\times$ Cyber-attacks on demand $\times$ Emergency MILP; CPLEX G
[80] $\times$ Deterministic line and pipeline outages $\times$ Restoration MISOCP; Relaxation inducement K
[74] Load demand Line outages Scenarios, ambiguity sets Restoration Chance-constrained adaptive distributionally robust MILP; CPLEX K
[57] $\times$ Deterministic large-scale blackouts $\times$ Restoration MILP L
[81] Wind forecast Deterministic line outages Scenarios Restoration Stochastic MILP; Gurobi I
[78] Renewable outputs, electric and thermal loads Deterministic line outages Scenarios Restoration Stochastic MILP; Gurobi J
[135] $\times$ Power line and pipeline outages $\times$ Restoration MILP, Gurobi M
@@
Figure 13: The framework of the two-level restoration strategy for a distribution network with HRSs[78]
After the extreme events, some measures related to system topologies or resources could be adopted to restore load. For example, Zhu et al.[80] developed a resilience-oriented service restoration strategy for a power-hydrogen distribution system by taking some measures, e.g., network topology reconfiguration, scheduling of dispatchable generators and HESSs, and load prioritization. Similarly, Afsari et al.[74] proposed a service restoration strategy for microgrids with hydrogen storage based on network topology reconfiguration and the coordination of HESSs and hydrogen-based micro-gas turbines. Before network reconfiguration is performed, an HMES black-start is required. In [57], Lu et al. developed a black start strategy for hydrogen-integrated renewable grids with energy storage systems, where hydrogen fuel cells are used as black start resources. In addition to stationary HERs, MHERs could also be scheduled for restoration response. For example, Su et al.[81] proposed a restoration strategy for hydrogen-accommodated microgrids based on a two-stage stochastic programming framework, where the location allocation of mobile wind turbines is decided in the first stage according to the shortest-path information in the transportation network considering uncertainties in wind generation. Then, the expected costs of power outages will be minimized by scheduling mobile wind turbines and HESSs in microgrids jointly. In [78], Zhao et al. proposed a two-level restoration strategy for a distribution network with HRSs. As shown in Fig. 13, the distribution system operator schedules mobile energy sources (i.e., HTTs and HFCVs) to minimize the power imbalance among microgrids and reduce the unserved load in the upper level. Then, based on the allocation results, a microgrid energy management strategy is designed to minimize the system operation cost of each microgrid. Simulation results showed that the coordination of mobile energy sources and HFCVs can reduce the power imbalance and improve the self-healing ability of hydrogen-electricity systems. Different from above works, Wang et al.[135] proposed a restoration strategy for an integrated power and hydrogen distribution system with the consideration of power line and hydrogen pipeline outages caused by extreme events. When power outages occur, reconfigurations of transmission line and hydrogen pipeline are first adopted to avoid islanding. Then, repair crews are dispatched to repair the fault lines and hydrogen pipelines. In addition, mobile battery-carried vehicles are dispatched to cooperate with repair crews to restore critical loads.
Figure 14: The framework of the rolling dispatch method with an alert mechanism[83]
@@ VII-B Multi-stage operation strategy
Although the above single-stage operation strategies can enhance HMES resilience, they cannot implement the coordination between different stages. To this end, some multi-stage operation strategies for HMESs have been proposed. For example, Haggi et al. proposed a proactive rolling horizon optimization-based method for hydrogen systems to enhance the power system resilience[50], which can implement the seamless switching between normal and emergency modes. Under normal mode, hydrogen systems assist the power grid by acting as a load. Once the extreme event time is known based on the forecast, the hydrogen tank is filled to make preparation for emergency mode. In [87], Liu et al. developed a risk-averse receding horizon optimization method to support the proactive and emergency operation of the industrial park based on integrated hydrogen-electricity-heat microgrids, where the measure of proactive operation includes increasing the reserve capacity of FCs and that of emergency operation is to reschedule components for the electricity and heat load survivability. In [83], Wang et al. proposed a rolling dispatch method to enhance the resilience of hydrogen-penetrated Antarctic energy systems under extreme weather. The framework of the rolling dispatch method can be observed in Fig. 14, where three stages could be identified, i.e., normal scheduling, alert scheduling, and emergency scheduling. Moreover, the above-mentioned stages will transfer from one to another according to different event states. To be specific, when the extreme event information is known based on forecasting techniques at slot $t+s$, the normal scheduling stage with horizon length $h$ will transfer to the alert scheduling stage with rolling horizon length $h+\tau$. Here, $\tau$ is the extreme event duration. When the event beginning is detected, the alert scheduling stage will transfer to the emergency scheduling stage with a rolling horizon length $\tau$. Once the event ending is detected, the alert scheduling stage will transfer to the normal scheduling stage. Since the function of alert scheduling is similar to that of preventive response (i.e., making preparation for emergency response), their roles are regarded as equivalent in Table XI. Similar rolling horizon optimization-based methods considering economic operation, proactive preparation, and emergency operation can be observed in [90].
TABLE XII: Summary of Operation Enhancement for HMES (Multi-stage)
@@
Ref. Parameter/event uncertainty Contingency uncertainty Uncertainty processing methods Operation stages Problem types and solving methods Hydrogen-related measure types
@@
[87] $\times$ NkN-k line contingencies Scenarios Preventive, emergency Stochastic MISOCP; Receding horizon method A, D
@@
[83] Renewable energy Outage of renewable energy units Box uncertainty set Preventive, emergency MILP; Gurobi A, D
@@
[90] Extreme event starting time Deterministic wind generation curtailment Scenarios Preventive, emergency Rolling-horizon MILP; Gurobi A, D
@@
[72] $\times$ Failed transmission lines, their failure, and repair times Scenarios Preventive, emergency Stochastic multi-objective MILP D
[23] $\times$ Component outages Scenarios Preventive, emergency MILP; Gurobi A, D
@@
[11] $\times$ Line faults Uncertainty sets Preventive, emergency Two stage robust optimization; C&CG B, F
[73] $\times$ Line outages Scenarios Preventive, emergency Stochastic MILP C, D
[77] $\times$ NkN-k line contingencies Uncertainty sets Preventive, emergency Tri-level MILP; Nested C&CG C, F
@@
[82] Load, renewable energy, and travelling time Deterministic grid outages Scenarios Preventive, restoration Stochastic MILP; Gurobi D
@@
[88] $\times$ Lines/component failure Scenarios Preventive, restoration Stochastic multi-objective programming; NSGA-II C, J
@@
[75] $\times$ Deterministic line outage $\times$ Preventive, restoration MISOCP; CPLEX A, J, K
[138] $\times$ Deterministic grid outage $\times$ Emergency, restoration MILP; CPLEX G, J
@@
[21] Renewable generation, load demands, EV charging level Deterministic line outage Scenarios Emergency, restoration Bi-level MILP; Adaptive ADMM algorithm G, J
@@
[139] $\times$ Deterministic line outages and insufficient gas supply $\times$ Emergency, restoration Stochastic MIQCP; Augmented ADMM algorithm F, D, J
@@
TABLE XIII: Representative Hydrogen-related Operation Measures for HMES resilience enhancement
@@
Type-A HESS pre-charging Increasing HESS level for reserve[50][75][83][85, 86, 87][23]
Preventive Type-B HTT scheduling HRS-based microgrid formation assisted by transport-based measures (e.g., HTTs and mobile hydrogen trucks) [30], pre-disaster HTT scheduling[11]
@@
Type-D HESS discharging Hydrogen tank-driven fuel cell for independent power supply during outages[51], discharging batteries and HESS[90], HESS and CHP work at maximum power[83], discharging HESS for serving HFCVs[23], discharging HESS for power supply[87]
Type-E Natural gas/hydrogen pipeline-assisted energy supply Utilizing hydrogen enriched natural gas for fuel support[31], using pipelines as large-scale energy storage[136], using hydrogen pipeline for power support[134]
Emergency Type-F HTT/hydrogen truck scheduling Using mobile hydrogen trucks for critical load support[18][19], fuel cell power supply supported by hydrogen from HTTs[11], dynamic re-routing of HTTs[77]
@@
Type-I Mobile hydrogen storage Combined operation of mobile wind turbines and HESSs[81]
Type-J HFCV scheduling HFCV routing and dispatch[75][78][88], HFCV for power support[75]
Restoration Type-K HESS scheduling Scheduling HESS with other components[82], HESS discharging[74][75]